OpenStudy (anonymous):

(a 3 - 2a + 5) - (4a 3 - 5a 2 + a - 2)

alones (alones):

Okay a(3)−2a+5−(4a(3)−5a(2)+a−2) distribute the negative sign: a(3)−2a+5+−1(4a(3)−5a(2)+a−2) a(3)+−2a+5+−1(4a(3))+−1(−10a)+−1a+(−1)(−2) a(3)+−2a+5+−12a+10a+−a+2 3a+−2a+5+−12a+10a+−a+2 then combine like terms: =3a+−2a+5+−12a+10a+−a+2 =(3a+−2a+−12a+10a+−a)+(5+2) =−2a+7

OpenStudy (ijlal):

Simplify it simply :D remember the negative sign times itself is positive example (-2)(-2)=4

OpenStudy (ijlal):